OpenStudy (anonymous):

How many square inches of tin are needed to make the top of a can that is 10 inches in diameter?

jimthompson5910 (jim_thompson5910):

hint: area of a circle is A = pi*r^2 where A = area r = radius pi = 3.14 (approximately)

jimthompson5910 (jim_thompson5910):

remember that d = 2*r where d = diameter r = radius

OpenStudy (mathmale):

So, Khmer, try putting this all together to determine the area of the can top if the diameter of the can top is 10 units.